OpenStudy (anonymous):

List x1, x2, x3, x4 where xi is the right endpoint of the four equal intervals used to estimate the area under the curve of f(x) between x = 3 and x = 5.

OpenStudy (zarkon):

\[x_i=3+i\times\Delta x\] where \(\Delta x=\dfrac{5-3}{4}\)

OpenStudy (anonymous):

Would I use 3 4 and 5 as my i values?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[\Delta x=\frac{5-3}{4}=\frac{2}{4}=\frac{1}{2}\] \[x_i=3+i\times\frac{1}{2}\] \[x_1=3+1\times\frac{1}{2}=3.5\] \[\cdots\]